1. Read the following job descriptions and decide on a percentage pay increase for each of the eight employees.

2. Make salary increase recommendations for each of the eight managers that you supervise. There are no formal company restrictions on the size of raises you give, but the total for everyone should not exceed the $10,900 (a 4% increase in the salary pool) that has based budgeted for this purpose. You have a variety of information on which to base the decisions, including a "productivity index" (PI), in which industrial engineering computes as a quantitative measure f operating efficiency for each managers work unit. This index ranges from a high 10 to a low 1. Indicate the percentage increase you would give each manager in the blank space next to the managers name. Be prepare to explain why?

___A. Alvarez. Alvarez is new this year and has a tough workgroup whose task is dirty and difficult. This is a hard position to fill, but you don't feel Alvarez is particularly good. The word around is that the other managers agree with you. PI = 3, Salary = $33,000

___B. J. Cook. Cook is single and a "swinger" who enjoys leisure time. Everyone laughs at the problems Cook has getting the work out, and you feel it certainly is lacking. Cook has been in the job two years. PI = 2, Salary = $34,500

___Z. Davis. In the position three years, Davis is one of your best people, even though some of the other managers don't agree. With a spouse who is independently wealthy, Davis doesn't need money but likes to work. PI = 7, Salary = $36,600

___M. Frame. Frame has personal problems and is hurting financially. Others gossip about Frame's performance, but you are quite satisfied with this second-year employee. PI = 7, Salary = $34,700

___C.M.Liu. Liu is just finish a fine first year in a tough job. Highly respected by the others, Liu has a job offer in another company at a 15% increase in salary. You are impressed, and the word is that the money is important. PI = 9, Salary = $34,000.

___B. Ratin. Ratin is a first-year manager whom you and the others think is doing a good job. This is a bit surprising since Ratin turned out to be a "free spirit" who doesn't seem to care much about money or status. PI = 9, Salary = $33,800.

___H. Smith. Smith is a first year manager recently divorced and with two children to support as a single parent. The others like Smith a lot, but your evaluation is not very high. Smith could certainly use extra money or statues. PI = 5, Salary = $33000.

___G. White. White is a big spender who always has the latest clothes and a new car. In the first year on what you would call an easy job, White doesn't seem to be doing very well. For some reason, though, the others talk about White as the "cream of the new crop," PI = 5, Salary = $33,000.

  • What was your overall strategy for determining your approach to the annual pay raise exercise?
  • Did you include all the information? Did you omit any information?
  • What were your challenges? What employee(s), if any, did you find most challenging when making your decisions?
  • Describe the equity theory of motivation. Why is it important?
Answer #1

The best approach here to be extremely objective in providing the salary increase. This means that we will create a weight based system for the allocation of the raise.

The total weight in terms of productivity index is 3+2+7+7+9+9+5+5 = 47.

A Alvarez = 10900*3/47 = 695.7

J Cook = 10900*2/47 = 463.8

Z. Davis = 10900*7/47 = 1623.4

M. Frame = 10900*7/47 = 1623.4

M. Liu = 10900*9/47 = 2087.2

B. Ratin = 10900*9/47 = 2087.2

H. Smith = 10900*5/47 = 1159.5

G. White = 10900*5/47 = 1159.5

The overall strategy for determining the annual pay rise exercise is to be completely objective and base the pay rise on the productivity index. This is a quantitative application and thus is the fairest method in order to allocate the funds. Also the difference in the base salary of these employees are mostly within the range of $3600 of each other and hence the displacement in terms of percentage will not be excessive.

I have included the PI information. All other background information has been omitted. This is because, irrespective of their personal life, we as an organization need to be professional and allocated fund based on their performance and not necessarily their personal story. This is why even though Davis does not need money and Frame needs the money, since both their performance is equal, they are given equal raise.

The key challenge is to remain fair and the dilemma to consider their personal information. For example, making the decision for M Liu was particularly hard. He has an offer for 15% hike at a competitor. However, with his current performance, I can at most give him a hike that corresponds to 6.1%. This poses a major risk of losing him. However, between the decision to be fair and to lose him, I chose to be fair.

Equity theory of motivation states that employees’ motivation is directly linked to their sense of equity and fairness. If the employees feel that the organization is unfair or unjust then they will be demotivated.